# Distributed AI Platform for the 6G RAN

### v1

@@ Abstract
Cellular Radio Access Networks (RANs) are rapidly evolving towards 6G, driven by the need to reduce costs and introduce new revenue streams for operators and enterprises. In this context, AI emerges as a key enabler in solving complex RAN problems spanning both the management and application domains. Unfortunately, and despite the undeniable promise of AI, several practical challenges still remain, hindering the widespread adoption of AI applications in the RAN space. This article attempts to shed light to these challenges and argues that existing approaches in addressing them are inadequate for realizing the vision of a truly AI-native 6G network. Motivated by this lack of solutions, it proposes a generic distributed AI platform architecture, tailored to the needs of an AI-native RAN and discusses its alignment with ongoing standardization efforts.
Index Terms:
6G RAN, Distributed AI, Radio Access Network, vRAN, AI RAN.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
I Introduction
Cellular Radio Access Networks (RANs) are undergoing a paradigm shift. The main driver behind this evolution is the reduction of the high CapEx and OpEx that telco operators are faced with and the introduction of new revenue streams. This transformation started with the launch of 5G, which introduced several key changes in the way that the network is being deployed and managed. As illustrated in Fig. 1, it turned the monolithic base stations into disaggregated and virtualized components – Central Unit (CU), Distributed Unit (DU) and Radio Unit (RU) – that can be deployed on top of commodity hardware across several locations (far edge, near edge, cloud), simplifying their lifecycle management and release of new features. Furthermore, it introduced open and programmable interfaces, allowing the deployment of applications on top of RAN Intelligent Controllers (RICs), to further accelerate innovation. Finally, it introduced new radio access technologies, like Massive MIMO and millimeter waves, and pushed the densification of the network to new limits, with the goal of expanding the network capacity and enabling new types of applications (e.g., IoT).
@@ I Introduction
Figure 1: High-level overview of the 5G RAN. The virtualized RAN infrastructure capabilities can vary depending on the location.
While the shift to 5G has so far served as a transitional phase, the emerging 5G Advanced and 6G networks are expected to truly unlock the networks’ potential in two ways. First, by developing sophisticated control and management solutions that can tackle a plethora of long-withstanding network problems (e.g., in the context of radio resource management, mobility, energy savings, etc.), which have been further amplified by the added complexity of 5G. Second, by introducing new transformative applications (e.g., joint communication-sensing, security, slicing), that can add value and provide differentiation for operators and verticals.
In this context, AI is expected to play a pivotal role. The recent advances in the domain of generative AI have acted as a catalyst in drawing the attention of the telecommunications industry to the benefits of AI in solving the aforementioned problems. AI is an ideal fit for many long standing RAN problems, which involve pattern recognition of signals for classification, predicting traffic and devising solutions for computationally intractable problems, such as scheduling. In fact, the influence of AI is becoming so big in the telco space, that the general consensus is that next generation mobile networks should be AI-native, with both the industry and academia rallying behind initiatives like the AI-RAN Alliance [1].
Unfortunately, while the promise of AI in the RAN is undoubtedly high, its potential still remains untapped, due to several practical roadblocks. First, the high accuracy of AI models is contingent to the existence of rich data sources. Given that RAN base stations are distributed across tens of thousands of locations, one would have to deal with several problems related to the collection and transmission of vast amounts of data. Second, RAN applications leveraging AI models present heterogeneous characteristics in terms of their compute requirements, response latency and privacy constraints. When operating over a distributed hierarchy of edges and the cloud, the varied compute availabilities and network bandwidths pose a significant challenge for operational deployments and handling AI RAN application constraints is far from straightforward.
Motivated by these observations, in this article, we attempt to shed light to the challenges of designing and deploying AI models for the RAN. We argue that existing approaches, which rely on the static data collection and orchestration of AI models are not suitable to meet the demands of AI-native 6G networks. Building on these observations, we present our vision of a distributed AI architecture that is tailored to the needs and requirements of an AI-native RAN, and we outline our thoughts of how our proposed architecture can be aligned with the ongoing standardization efforts.
II The Need for AI in RAN
Here, we discuss why the need for AI in the RAN space is becoming increasingly important. The AI-RAN Alliance has succinctly grouped AI use cases in three key domains [1]:
1) AI-for-RAN – This refers to utilizing AI to optimize and enhance the RAN performance, motivated by the need to use the limited spectrum efficiently. The demand for more traffic has led to the introduction of new, higher frequency bands. These bands offer more capacity but lower coverage, leading to an increase in the overall tower density. Furthermore, new radio resource allocation technologies, like Massive MIMO, see the use of large antenna arrays as a way to further increase the cell capacity. Assigning users to different bands and managing interference across cells by adjusting scheduling, power control and antenna assignments leads to a search space explosion, which is computationally intractable to manage with conventional optimization methods, making the use of AI a compelling alternative [2].
Similar to radio resource optimization, there exist a whole set of infrastructure optimization problems, caused by the increased 5G complexity, which AI promises to solve. Telco operators often manage close to 100k RAN sites. Predictive maintenance for an infrastructure at that scale (e.g., dealing with hardware issues, software bugs, etc.) is a challenge on its own. The virtualization of the RAN adds an extra layer of complexity, making the identification and root cause analysis of performance issues far from obvious [3]. Finally, in the context of energy efficiency, a lot of effort has been on how to reduce the power consumption of various elements, while maintaining the network coverage and quality, with several AI-based research works showing promising results [4].
@@ II The Need for AI in RAN
In contrast to cloud computing, where the sharing of compute resources is commonplace, sharing resources with RAN workloads introduces significant challenges. The real-time nature of the RAN means that sharing can lead to deadline violations and adversely affect network performance. Several recent proposals have demonstrated how AI can be leveraged to share the infrastructure in both CPUs [6] and GPUs [7].
3) AI-on-RAN – This domain encompasses use cases that leverage the RAN infrastructure to support AI applications. A key enabler of such use cases are the open interfaces exposed by the RAN functions, which can allow third-parties to tap into RAN data and/or affect RAN traffic-related decisions, in order to enhance the application capabilities. One example that falls under this category is localization via sensing. By leveraging channel state information exposed by the physical layer of the RAN, one can perform user localization, which can be important for several use cases, like asset tracking and self-driving cars [8]. AI-based techniques leveraging physical layer information can also be employed to provide context-aware security [9], in consideration of battery-constrained devices. Finally, more traditional AI-based applications, like video analytics could also leverage the interplay with the RAN, to accommodate latency critical use-cases (e.g., [10]).
III Background
III-A Canonical application examples
To help with highlighting the need for a distributed AI RAN platform, we define two AI application examples (Fig. 2), which we use as a reference for the remainder of this article.
Figure 2: Graphs of AI/ML models for RAN applications
RAN slicing scheduler – A scheduler allocates radio resources across network slices (inter-slice scheduling) and across users of the same slice (intra-slice scheduling). Both schedulers use the traffic demand information provided by the phones through buffer status reports, to try and predict the traffic load for the upcoming period [11]. They also use physical layer sounding reference signals to predict the users’ signal quality. The inter-slice scheduler makes scheduling decisions at coarse granularities (e.g., seconds) and feeds its decisions to the real-time intra-slice scheduler. This example illustrates both AI-on-RAN (predicting load using buffer status reports) as well as AI-for-RAN (scheduling decisions).
Anomaly detection and root cause analysis – A service management and orchestration framework tries to detect RAN-related anomalies and identify their root cause for potential mitigations towards AI-for-RAN. Since the anomalies can originate both from the RAN and the platform, the application collects data from both sources [3]. It also applies imputation and aggregation techniques, to deal with noisy/missing data and to reduce their volume. It then processes the collected statistics, to detect if an anomaly is present, and if so, it attempts to localize it and detect its root cause.
From these examples, we see that a typical AI RAN application, rather than being a monolith, may consist of a sequence of blocks. Each block can do its own data pre-processing and can include an ML model in itself, with substantial research and development driving its release. Those blocks are chained together into a graph (e.g., by solution providers), where the output of one model (e.g., mobility predictor) becomes the input of the next (e.g., resource block scheduler). Typically the traffic volume between components reduces from left to right, as more data gets processed and aggregated (signified with narrower arrows in the examples).
III-B Distributed edge infrastructure
IV Challenges to build AI RAN applications
IV-A Data collection
Different applications require different feature sets, that might have heterogeneous characteristics in terms of type, time granularity, etc. For example, the radio resource scheduling application might require real-time data from the RAN network functions, while the anomaly detection might need to combine aggregate data from both the RAN and the platform in time windows of seconds. Similarly, the anomaly detection application might rely on the inter-arrival time between IQ samples, while the scheduler might require the actual raw IQ samples carrying the sounding reference signals.
The current standard practice to bypass this roadblock is to expose a set of coarse-grained data sources, that are applicable to a wide range of use cases. These data sources are exposed through static APIs specified by standardization bodies like 3GPP and O-RAN. For instance, O-RAN defines the E2 interface for the collection of pre-defined RAN KPIs in the form of so-called static “service models”, and the O2 interface for the collection of monitoring data from the platform.Any change to a data source (e.g. reporting the 90th percentile of packet latency instead of max), requires the standardization body consensus.This can be a long process that is typically met with skepticism, due to performance concerns that might arise when modifying mission critical RAN and platform code.
In conclusion, data collection challenges force AI RAN applications to be developed and deployed as an afterthought, subject to the available data sources, rather than in a truly AI-native way, where the data collection is application-driven.
IV-B RAN AI application orchestration
The disaggregated nature of the RAN means that its network functions might be deployed across the edges and the cloud. This raises a fundamental question as to what should be the location in which the blocks of AI applications should reside. Answering this question is far from straightforward. It involves matching the applications’ requirements to the capabilities of the infrastructure, which can vary in terms of compute resources, network bandwidth, latency, etc, depending on the location (Fig. 1). The placement decision could also be affected by other factors, such as privacy concerns, since operators might have a preference or legal obligation to process certain data streams in their own premises.
The many constraints and trade-offs make the deployment of AI RAN applications a major challenge. First, a developer needs to understand the underlying constraints which may differ from one deployment to another. Second, they need to carefully choose what data to collect and where to place the various application blocks, to maintain high accuracy, while respecting the constraints. Third, deployed applications will naturally have competing requests for the same resources, meaning that it falls on the developer to design their application with placement flexibility in mind. Finally, as alluded to in the case of AI-and-RAN, AI RAN applications have to coexist with other AI applications (non-RAN) that also execute on the edge such as video analytics and self-driving cars [10]. In the absence of a structured framework, AI RAN applications have to be developed with ad hoc programming interfaces and manually distributed by developers. As a result, they are rarely deployed at scale in production settings and have led to their potential remaining largely untapped.
V Distributed AI platform for RAN
Figure 3: Distributed AI platform architecture.
While this approach provides flexibility in tailoring the data collection process to the AI application’s requirements, it also introduces safety concerns (e.g., illegal memory accesses). For this, we propose to use probes based on the eBPF technology, which has recently caught the attention of the telco industry [12, 13]. eBPF allows the injection of code to instrumentation points, subject to a static verification process that guarantees the safety of the injected code. While eBPF originates in the Linux kernel, recent advances [12] have expanded its use to userspace RAN applications.
AI processor runtime – Considering the constraints and trade-offs that characterize the various locations of the infrastructure, an AI-native platform should allow the seamless deployment of AI applications to the most suitable location, without the developer having to worry about providing location-specific flavors (i.e., implementations). In the proposed architecture, this is achieved through the AI processor runtime. The runtime can be deployed at any location where AI applications are expected to run and introduces an API for its interactions with the applications. The runtime should provide the following functionalities:
– Data exchange: It should allow the exchange of data streams amongst the blocks of AI applications deployed in different locations through a common message bus. Similarly, it should enable the exposure and issuing of RPCs for applying control decisions.
– Execution environment: It should implement the process of running inference or training tasks, by abstracting the underlying compute resources (e.g., CPUs and GPUs) and exposing them to the AI applications, catering to both RAN and non-RAN AI applications.
The distributed AI RAN platform does not prescribe an implementation for the AI processor runtime, as long as it provides the functionalities described above. We eschew making a prescription because we believe that the framework should be extensible to include existing and future runtime environments and messaging technologies, thus not restraining AI developers. As an example, one could consider using Docker containers combined with a WebAssembly (WASM) runtime as a highly portable and sandboxed execution environment, while the message bus implementation could be based on REST or gRPC calls.
Applications are exposed to the orchestrator in the form of blocks in a graph.The blocks are characterized by requirements in terms of compute, memory, network, latency, as well as other constraints (e.g., the locations where the block is not allowed to run due to privacy concerns). The orchestrator takes into account both the developer requirements and the capabilities of the infrastructure and places the blocks to the AI processor runtimes that maximize the overall utility of the platform. Unique to AI workloads are inference parameters that influence resource demand [10], and hence is a resource allocation knob for the orchestrator. Examples of inference parameters would be data sampling rates, or AI models for anomaly detection. The sampling rate directly impacts the compute and network demand, while the choice of model dicates the compute demand, including the necessity for GPUs. We propose the AI RAN applications to expose their inference parameters to the orchestrator, and design the orchestrator to dynamically change the inference parameters.
Figure 4: Distributed deployment scenarios for orchestrated AI applications.
VI An efficient AI processor runtime for the far edge
We advocate that a highly optimized far edge runtime is required, as it is expected to host real-time AI applications (e.g. the RAN slicing scheduler), as envisaged with a concept of dApps [14]. As such, it needs to have sub-millisecond reaction times. Furthermore, the far edge is resource constrained, with only a small fraction of CPUs available for AI applications, and with a small or no GPU present. Our proposed design is illustrated in Fig. 5 and has the following characteristics:
Figure 5: Architecture of far edge AI processor runtime.
Tight integration with the probes –
Efficient, flexible and secure execution environment –
Admission control & real-time resource allocation –
VII Open vs Closed architecture and interfaces for integration with RAN

### v2

@@ Abstract.
Cellular Radio Access Networks (RANs) are rapidly evolving towards 6G, driven by the need to reduce costs and introduce new revenue streams for operators and enterprises. In this context, AI emerges as a key enabler in solving complex RAN problems spanning both the management and application domains. Unfortunately, and despite the undeniable promise of AI, several practical challenges still remain, hindering the widespread adoption of AI applications in the RAN space. In this work, we attempt to shed light to these challenges and argue that existing approaches in addressing them are inadequate for realizing the vision of a truly AI-native 6G network. We propose a distributed AI platform architecture, tailored to the needs of an AI-native RAN.
copyright: none
1. Introduction
In 5G, RAN is undergoing a paradigm shift, disaggregating monolithic base stations into separate components (CU, DU, and RU) and containerizing them on commodity hardware/software platforms across edges and cloud (Fig. 1). This enables open interfaces for RIC-based applications, accelerating innovation and facilitating new capacity-boosting technologies like Massive MIMO, mmWaves, and densification. 6G will further address long-standing problems (e.g., radio resource management, mobility, energy savings) and introduce transformative uses (e.g., joint communication-sensing, security, slicing).
The AI revolution, as catalyzed by the current generative AI phenomenon, has motivated the telecom industry to see AI as an ideal fit for RAN problems like signal classification, traffic prediction, and optimal scheduling. The next-gen mobile networks, including 6G, will be “AI-native”. AI-RAN Alliance (Alliance, 2024), a telecom industry’s AI initiative, has defined three promising domains of AI use cases in the RAN:
(1) AI-for-RAN – use AI to optimize RAN performance. For example, higher frequencies, cell density, and larger antenna arrays will increase the search space in scheduling with AI-based solutions (Bartsiokas et al., 2022). Infrastructure management like predictive maintenance of RAN sites, vRAN troubleshooting (Sun et al., 2024), and energy savings (Kundu et al., 2024; Singh et al., 2021; Kalia et al., 2025) will all benefit from AI.
(2) AI-and-RAN – share compute (CPU/GPU) between RAN and AI apps. The vRAN is largely underutilized ($50\%$(Nvidia, 2023) due to overprovisioning and traffic inbalance (Foukas and Radunovic, 2021), but AI can be leveraged to achieve sharing without violating the RAN’s realtimeness (Foukas and Radunovic, 2021; Schiavo et al., 2024; Foukas and Radunovic, 2025).
(3) AI-on-RAN – use RAN infrastructure for AI apps. For example, localization from channel data for tracking/autonomy (Trevlakis et al., 2023), context-aware security (Chorti et al., 2022), and latency-critical video analytics (Zhang et al., 2024) will benefit from interfacing with RAN.
Despite these promises, challenges persist. First, we need to collect data from vastly and geo-distributed RAN sites for accurate AI models. Second, AI-RAN use cases vary in compute, latency, and privacy needs, complicating edge-cloud orchestration. The goal of this research is to design a distributed AI platform for 6G RAN. We highlight the issues, arguing static approaches fall short for AI-native networks. We propose a distributed AI architecture to address these problems.
@@ 1. Introduction
Figure 1. High-level 5G RAN overview. The RAN infrastructure capabilities can vary depending on the location.
2. Background
2.1. Canonical application examples
To help with highlighting the need for a distributed AI RAN platform, we define two examples (Fig. 2), which we use as a reference for the remainder of this work.
RAN slicing scheduler – A scheduler allocates radio resources across network slices (inter-slice scheduling) and across users of the same slice (intra-slice scheduling). Both schedulers use the traffic demand information provided by the phones through buffer status reports, to try and predict the traffic load for the upcoming period (Balasingam et al., 2024; Foukas et al., 2017). They also use physical layer sounding reference signals to predict the users’ signal quality. The inter-slice scheduler makes scheduling decisions at coarse granularities (e.g., seconds) and feeds its decisions to the real-time intra-slice scheduler. This example illustrates both AI-on-RAN (predicting load using buffer status reports) as well as AI-for-RAN (scheduling decisions).
Figure 2. Graphs of AI/ML models for RAN applications
Anomaly detection and root cause analysis – A service management and orchestration framework tries to detect RAN-related anomalies and identify their root cause for potential mitigations towards AI-for-RAN. Since the anomalies can originate both from the RAN and the platform, the application collects data from both sources (Sun et al., 2024). It also applies imputation and aggregation techniques, to deal with noisy/missing data and to reduce their volume. It then processes the collected statistics, to detect if an anomaly is present, and if so, it attempts to localize it and detect its root cause.
From these examples, we see that a typical AI RAN application, rather than being a monolith, may consist of a sequence of blocks. Each block can do its own data pre-processing and can include an ML model. Those blocks are chained together into a graph, where the output of one model becomes the input to the next. Typically the traffic volume between components reduces from left to right, as more data gets processed and aggregated.
2.2. Distributed edge infrastructure
3. Challenges to build AI RAN applications
3.1. Data collection
Different applications require different feature sets, that might have heterogeneous characteristics in terms of type, time granularity, etc. For example, the radio resource scheduling application might require real-time data from the RAN network functions, while the anomaly detection might need to combine aggregate data from both the RAN and the platform in a time windows of seconds. Similarly, the anomaly detection application might rely on the inter-arrival time between IQ samples, while the scheduler might require the actual raw IQ samples carrying the sounding reference signals.
The current standard practice to bypass this roadblock is to expose a set of coarse-grained data sources, that are applicable to a wide range of use cases. These data sources are exposed through static APIs specified by standardization bodies like 3GPP and O-RAN. For instance, O-RAN defines the E2 interface for the collection of pre-defined RAN KPIs, and the O2 interface for the collection of platform data.Any change to a data source requires the standardization body consensus, which can be a very long process. As such, AI RAN applications end up being developed as an afterthought, subject to the available data sources, rather than in a truly AI-native way, where the data collection is application-driven.
3.2. RAN AI application orchestration
The disaggregated nature of the RAN means that its network functions might be deployed across the edges and the cloud. This raises a fundamental question as to what should be the location in which the blocks of AI applications should reside. Answering this question is far from straightforward. It involves matching the applications’ requirements to the capabilities of the infrastructure, which can vary in terms of compute resources, network bandwidth, latency (Fig. 1) and privacy.
The many constraints and trade-offs make the deployment of AI RAN applications a major challenge. First, a developer needs to understand the underlying constraints which may differ from one deployment to another. Second, they need to carefully choose what data to collect and where to place the application blocks, to maintain high accuracy, while respecting the constraints. Third, deployed applications will naturally have competing requests for the same resources, meaning that it falls on the developer to design their application with placement flexibility in mind. Finally, AI RAN applications have to coexist with other AI applications that also execute on the edge, such as video analytics and self-driving cars (Zhang et al., 2024). In the absence of a structured framework, AI RAN applications have to be developed with ad hoc programming interfaces and manually distributed by developers.
4. Distributed AI platform for RAN
Figure 3. Distributed AI platform architecture.
While this approach provides flexibility in tailoring the data collection process to the AI application’s requirements, it also introduces safety concerns (e.g., illegal memory accesses). For this, we propose to use probes based on the eBPF technology, which has recently caught the attention of the telco industry (Foukas et al., 2023; Soldani et al., 2023). eBPF allows the injection of code to instrumentation points, subject to a static verification process that guarantees the safety of the injected code. While eBPF originates in the Linux kernel, recent advances (Foukas et al., 2023) have expanded its use to userspace RAN applications.
AI processor runtime – Considering the constraints and trade-offs that characterize the various locations of the infrastructure, an AI-native platform should allow the seamless deployment of AI applications to the most suitable location, without the developer having to worry about providing location-specific flavors. In the proposed architecture, this is achieved through the AI processor runtime. The runtime can be deployed at any location where AI applications are expected to run and introduces an API for its interactions with the applications. The runtime should provide the following functionalities:
– Data exchange: It should allow the exchange of data streams among the AI application blocks deployed in different locations through a common message bus. Similarly, it should enable the issuing of RPCs for applying control decisions.
– Execution environment: It should implement the process of running inference or training tasks, by abstracting the underlying compute resources (e.g., CPUs and GPUs) and exposing them to both RAN and non-RAN AI applications.
As long as it provides the functionalities described above, we eschew making an implementation prescription because we believe that the framework should be extensible to include existing and future runtime environments and messaging technologies, thus not restraining AI developers. As an example, one could consider using Docker containers combined with a WebAssembly (WASM) runtime as a highly portable and sandboxed execution environment, while the message bus implementation could be based on REST or gRPC calls.
Figure 4. Distributed deployment scenarios for orchestrated AI applications.
Applications are exposed to the orchestrator in the form of blocks in a graph. The blocks are characterized compute, memory, network and latency requirements, as well as other constraints (e.g., the locations where the block is not allowed to run due to privacy concerns). The orchestrator takes into account both the developer requirements and the capabilities of the infrastructure and places the blocks to the AI processor runtimes that maximize the overall utility of the platform. Unique to AI workloads are inference parameters that influence resource demand (Zhang et al., 2024), and hence is a resource allocation knob for the orchestrator. Examples of inference parameters would be data sampling rates, or AI models with varying accuracy for anomaly detection. We propose the AI RAN applications to expose their parameters to the orchestrator, and design the orchestrator to dynamically change them.
5. An efficient far edge AI processor runtime
We advocate that a highly optimized far edge runtime is required to host real-time AI applications (e.g. the RAN slicing scheduler), as has been demonstrated in recent research works (e.g., (Ko et al., 2024)), and as envisaged in the O-RAN concept of dApps (D’Oro et al., 2022). As such, it needs to have sub-millisecond reaction times. Furthermore, the far edge is resource constrained, with only a small fraction of CPUs available for AI applications, and with a small or no GPU present. Our proposed design is illustrated in Fig. 5 and has the following characteristics:
Figure 5. Architecture of far edge AI processor runtime.
Tight integration with the probes –:
Efficient, flexible and secure execution environment – :
Admission control & real-time resource allocation – :
6. Implementation & Evaluation
To demonstrate the feasibility and benefits of our proposed architecture, we implemented a reference far edge AI processor runtime, based on the design outlined in Section 5, and also prototyped the distributed AI orchestrator.
6.1. Far edge AI processor runtime
Our reference AI processor runtime is written in $\sim$5K lines of C/C++ code and is publicly available at (Microsoft, 2025a). The runtime provides both a C++ and a Python API for the development of applications, to cater the needs of developers requiring either high performance (C++) or quick integration with popular AI frameworks and libraries (Python). It should be noted that the current reference implementation does not support WASM for deploying applications. The applications can instead be loaded in the form of shared libraries (.so), dynamically linked at runtime through a REST-based API. The extension to WASM applications is left as future work.
For the programmable in-kernel probes, we leveraged the popular eBPF technology that is part of the Linux kernel (eBPF, [n. d.]), while for the userspace probes, we leveraged the jbpf userspace eBPF framework (Microsoft, [n. d.]; Foukas et al., 2023), which has been designed with lightweight RAN instrumentation in mind. The integration of the probes with the AI processor runtime is done using zero-copy shared memory channels that are part of the jbpf framework.
@@ 6.1. Far edge AI processor runtime
Figure 6. Communication latency overhead for monitoring and control data to/from the AI processor runtime.
6.2. Distributed AI orchestrator
To demonstrate the benefits of distributed AI orchestration, we developed a proof of concept orchestrator. The orchestrator is capable of dynamically orchestrating container-based applications across the edge and the cloud, taking into account their GPU and network requirements.
For our evaluation, we consider a setup composed of a far edge location with a single HPE DL110 Linux server equipped with an Nvidia L4 GPU, with 24GB of memory, and an Azure cloud VM equipped with an Nvidia A100 GPU with 40GB of memory. As sample workloads we use the SpotLight (Sun et al., 2024) AI-based RAN anomlay detection and the visual SLAM of the Nvidia Isaac ROS framework (Nvidia, [n. d.]). The two applications were chosen as representative examples of AI-for-RAN and AI-on-RAN scenarios, and they both require a GPU.
[…]
As illustrated in Figure 7, we deploy SpotLight across 10 cells (case 1). The AI orchestrator decides to orchestrate all the SpotLight instances at the edge location, considering that the resources of the L4 GPU are adequate (utilization of $\sim$40%) and this deployment choice requires shipping no data to the cloud. Then, we request to deploy three instances of the visual SLAM application. The remaining edge GPU capacity is not adequate to accommodate the new workload, which requires 24GB of memory. As such, the AI orchestrator migrates the SpotLight applications to the cloud and deploys the visual SLAM applications to the edge (case 2). The reason behind this deployment choice of the orchestrator is that it maximizes the utilization of the edge GPU (100% utilization), while minimizing the amount of data that has to be shipped to the cloud. An alternative sub-optimal deployment option would be to deploy the visual SLAM to the cloud and keep SpotLight to the edge, however, this would lead to a requirement of shipping 120Mbps worth of data to the cloud (Case 2 - Alt). Along the lines of the discussion of Section 4, the orchestrator could also be tailored to take into account other optimization parameters, such as the latency or the deployment cost.
Figure 7. AI orchestrator deployment choices vs sub-optimal option.
7. Open vs Closed architecture and interfaces for integration with RAN